# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ 2.2 Enhanced Pop II star-formation efficiency
Small halos of mass $\sim 10^{5-9}\,\mathrm{M}_{\odot}$ are sensitive to the threshold for star formation $M_{\text{crit}}$, affected by feedback from streaming velocities, Lyman-Werner photons, and UV photons. These effects are incorporated using the thresholds:
$M_{\text{mol,crit}}\approx f_{v_{\text{bc}}}f_{\text{LW}}\left(5.8\times 10^{5}\,\mathrm{M}_{\odot}\right)\left(\dfrac{V_{c}}{4.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{3}\left(\dfrac{1+z}{21}\right)^{-3/2}$ (3)
$M_{\text{atm,crit}}\approx 10^{8}\,\mathrm{M}_{\odot}\left(\dfrac{V_{c,\text{atm}}}{17\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{3}\left(\dfrac{1+z}{10}\right)^{-3/2}$ (4)
where $f_{v_{\text{bc}}}=(1+v_{\text{bc}}/v_{\text{rms}})^{1.8}$ (Fialkov et al., 2012; Gessey-Jones et al., 2023) and $f_{\text{LW}}=(1+2J_{\text{21}})^{0.6}$ (Fialkov et al., 2013; Muñoz et al., 2022) account for the streaming velocity and Lyman-Werner feedback respectively, $V_{c}$ is the minimum circular velocity of halo for efficient cooling of gas, [^5] Note that the minimum circular virial velocity, defined as $V_{c}=\sqrt{GM_{\text{vir}}/R_{\text{vir}}}$, is simply related to a halo’s virial temperature often used in literature via: $T_{\text{vir}}=\dfrac{\mu m_{\text{p}}V_{c}^{2}}{2k_{\text{B}}}\approx 1300\,\text{K}\left(\dfrac{V_{c}}{4.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{2}$ (5) where $\mu=1.22$ is the mean molecular weight of neutral primordial gas, $m_{\text{p}}$ is the proton mass, and $k_{\text{B}}$ is the Boltzmann constant. and $V_{c,\text{atm}}=16.5\,\mathrm{km}\,\mathrm{s}^{-1}$. Finally,
$M_{\text{crit}}(\mathbf{x})=\begin{cases}\min(M_{\text{mol,crit}},M_{\text{atm,crit}})&\text{if }V_{c}<V_{c,\text{atm}}\\ M_{\text{mol,crit}}&\text{if }V_{c}\geq V_{c,\text{atm}}\end{cases}$ (6)
corresponding to molecular cooling and atomic cooling halos respectively. During the EoR, UV photons irradiate star-forming gas in ionized bubbles, further increasing $M_{\text{crit}}$ via photoheating feedback (described in Sobacchi & Mesinger, 2013; Cohen et al., 2016, accounting for ionized fractions in partially ionized cells).
@@
$f_{\star,\mathrm{II}}(\mathbf{x},M_{h},z)=\tilde{f}_{\star,\mathrm{II}}\begin{cases}0&M_{h}<M_{\text{crit}}\\ \dfrac{\log\left(M_{h}/M_{\text{crit}}\right)}{\log\left(M_{\text{atm}}/M_{\text{crit}}\right)}&M_{\text{crit}}\leq M_{h}<M_{\text{atm}}\\ 1&M_{\text{atm}}\leq M_{h}<M_{\text{high}}\\ \left(\dfrac{M_{h}}{M_{\text{high}}}\right)^{\alpha_{\star}}&M_{h}\geq M_{\text{high}}\end{cases}$ (7)
@@
and $f_{\star,\mathrm{II}}$ is naturally capped at $1$. The particular form of $M_{\text{high}}$ anchors the SFE at $z_{\text{min}}=6$, where our simulations end and the UVLF data is strongest (i.e., essentially calibrating the SFE magnitude and evolution to be with respect to redshift $z=6$). Figure 1 shows a schematic of this SFE model. Similar models with a flattening below a given mass threshold have been explored in literature before (Sun & Furlanetto, 2016; Sipple & Lidz, 2024), the latter of whom find that a shallow-slope best fits the lensed UVLFs from HST at low halo masses, as opposed to a single power law (dash-dotted green in Figure 1) that assumes a steep decline. The parameterization used in our work, due to its flexibility, relaxes the assumption of declining SFE made by the single power law model for small mass halos. We also expect our model to be reasonable compared to the double power law model (solid pink line in Figure 1) for two reasons: (i) AGN feedback effects are generally expected for large halos $\sim 10^{12}\,\mathrm{M}_{\odot}$. However, few halos reach $\sim 10^{11}\,\mathrm{M}_{\odot}$ at $z\geq 6$ in our simulation volume of interest $\sim 10^{8}\,\mathrm{cMpc}^{3}$. For comparison, multi-field determinations of UVLFs combining many HST/JWST surveys typically probe volumes of $\lesssim 10^{6}\,\mathrm{cMpc}^{3}$ across redshift bins, making large halos even rarer. Since the 21-cm signal, and in general reionization, is sensitive to the cumulative effect of all halos down to $10^{5}\,\mathrm{M}_{\odot}$, our relevant halo mass range is wider than other abundance matching and analytic inferences that usually consider $M_{h}\gtrsim 10^{9}\,\mathrm{M}_{\odot}$. (ii) The role of AGN at high redshifts is still quite uncertain. Although pre-JWST studies suggest a decline in AGNs at $z\gtrsim 6$ (Kulkarni et al., 2019), more recent works indicate a non-negligible faint AGN population (Harikane et al., 2023b; Fujimoto et al., 2024). The contribution of these AGN to star-formation quenching remains unclear since the bright end of the luminosity function is poorly sampled at high redshifts.
The redshift evolution factor $\beta_{\star}$ is a timely addition, motivated by the abundance of bright galaxies in recent high-redshift JWST observations. Physically, the term captures an an enhancement in the SFE in the early Universe, meaning DM halos of the same mass host more massive and brighter galaxies in the early Universe than today.
@@ 2.2.1 Phenomenological comment on the SFE
The SFE defined here is the fraction of the accreted gas in a dark matter halo that is converted into stars, sometimes referred to as the integrated baryon conversion efficiency (e.g. Moster et al., 2018; Zhang et al., 2022). To the first order, the SFE is equal to stellar mass – halo mass fraction up to a factor of the baryon fraction $f_{\text{b}}\approx 0.16$.
We note here that the SFE implementation in our sub-grid model is essentially an ensemble average since we do not track individual halos. For a given halo mass and redshift, the SFE of the galaxy in the host halo is fixed (i.e., there is no galaxy-to-galaxy scatter around the mean). This approach is reasonable as both the 21-cm signal and UVLF are statistical quantities probed at large scales (but see also Nikolić et al., 2024, on the importance of stochasticity). Observations of individual galaxies can significantly differ from the ‘global mean’, and indeed any observational dataset is biased to the brightest and most massive galaxies which will be inherently stochastic (e.g Mirocha & Furlanetto, 2023). Nonetheless, rare galaxies manifest in our simulations via our pixel-level formalism. The abundance of halos and their gas-to-baryon mass fraction ($f_{\text{g}}/f_{\text{b}}$) is conditional on the halo mass, and pixel-level overdensity and streaming velocity. The minimum mass for star-forming halos $M_{\text{crit}}$ is further affected by local feedback mechanisms. Thus, the stellar mass and UVLF still varies between pixels across the simulation. Rare overdensities at high-$z$ will produce rare high mass galaxies, which can have an elevated SFE if they fall in the power-law regime of Equation 7. There are two interpretations of enhancement at increasing redshifts:
@@
    The first interpretation is one of bursty star-formation, very much akin to “feedback-free bursts" (FFB), recently discussed in Dekel et al. (2023) and Li et al. (2024). Under this interpretation, the high SFE is a result of averaging over the star-formation history of a bursty galaxy with a short duty cycle. For example, if a galaxy has a burst of star-formation with $f_{\star,\mathrm{II}}\sim 1$ for $2\,\mathrm{Myr}$ followed by quiet phase of $f_{\star,\mathrm{II}}\sim 0$ for $8\,\mathrm{Myr}$, it corresponds to an overall SFE of $f_{\star,\mathrm{II}}\sim 0.2$ over the complete $10\,\mathrm{Myr}$ cycle. Hence, the SFE is the ratio of the bursty phase to the typical duration of each star-formation cycle, and multiple cycles/generations of starbursts occur in a single half-crossing time of a virialized DM halo. The mechanism for this burstiness, as Dekel et al. (2023) suggest, is rapid infall of gas via cold streams in low metallicity, high density environments. If this happens quicker than SNe/wind timescales, the gas collapse is essentially feedback-free. The halo mass threshold above which these bursty galaxies form decreases with increasing redshift in this scenario, leading to an early time enchancement (see Figure 2 of Li et al., 2024). Earlier work by Furlanetto & Mirocha (2022) on bursty star-formation during Cosmic Dawn also explores a boost in the SFE in small halos, although manifesting as a flattening of the SFE slope at the low mass end (see their Figure 5).
@@
    The second interpretation is the standard quiescent star-formation, through continuous gas supply to the galaxy via gradual accretion over the halo half-crossing time, but with a higher efficiency in the early Universe. This means that the star-formation history of the galaxy is uniform, and it is able to support a high SFE over long averaging time scales (say, $f_{\star,\mathrm{II}}\sim 0.2$ for $10\,\mathrm{Myr}$ as in the last example). The increased stellar mass fraction at high-$z$ could be due to a variety of reasons: a low feedback environment in the pre-reionization epoch (Barkana & Loeb, 2000); an increased binding energy / deeper potential wells of small halos giving rise to compact galaxies resilient to supernovae feedback (Dayal et al., 2014; Furlanetto et al., 2017); an increased penetration factor of gas into galaxies or decreased mass-loading factor of galactic outflows (Dekel & Mandelker, 2014); or positive feedback from SMBH-driven AGN (Silk et al., 2024).
@@ 2.3.1 The 21-cm signal
where the factor of $n_{0}$ and $n_{1}$ are number densities of hydrogen atoms in the lower and higher energy states, the factor of $3$ accounts for statistical degeneracy of the higher energy triplet state, ${\nu_{21}=1420\,\mathrm{MHz}}$ is the rest-frame frequency of the 21-cm line, and $T_{\mathrm{S}}$ is the spin temperature of the hydrogen atoms.
As the Universe expands, this 21-cm signal is observed today as a distortion to the CMB black body spectrum (or radio background, with temperature $T_{\text{r}}$) across the radio frequency range $\nu_{21,\text{obs}}=\nu_{21}/(1+z)$, corresponding to the different epochs at redshifts $z$. The strength of the signal is usually quantified in terms of the differential brightness temperature
$T_{21}(z)=\left(1-e^{-\tau_{21}}\right)\dfrac{T_{\mathrm{S}}-T_{\text{r}}}{1+z},$ (10)
@@
$\tau_{21}(z)\approx\dfrac{3}{32\pi}\dfrac{hc^{3}A_{10}}{k_{\text{B}}\nu_{21}^{2}}\dfrac{x_{\mathrm{HI}}(z)n_{\text{H}}(z)}{(1+z)dv_{\parallel}/dr_{\parallel}}\dfrac{1}{T_{\mathrm{S}}(z)}.$ (11)
with $x_{\text{HI}}$ being the fractional abundance of neutral hydrogen, $n_{\text{H}}$ the hydrogen number density, $A_{10}$ the spontaneous emission rate of 21-cm transition, and $dv_{\parallel}/dr_{\parallel}$ the proper velocity gradient along the observer line of sight. If ${T_{\text{S}}>T_{\text{r}}}$, the signal is seen in net emission against the radio background, otherwise it is seen in net absorption.
In order to calculate this brightness temperature, one needs the spatially and temporally varying fractional abundance $x_{\mathrm{HI}}$ and the spin temperature $T_{\mathrm{S}}$. The former depends on the distribution of ionizing stellar/quasar sources, while the latter is an interplay between several competing influences that can cause a spin-flip transition in hydrogen. Coupling of $T_{\mathrm{S}}$ to the background radiation temperature $T_{\text{r}}$ is mediated via scattering of background photons, while coupling to the kinetic temperature of matter $T_{\text{K}}$ is through collisions between atoms/electrons in dense IGM, and the Wouthusen-Field (WF) effect when star-formation begins (Wouthuysen, 1952; Field, 1958). Thus, we get the equation (Furlanetto et al., 2006; Venumadhav et al., 2018):
$T_{\mathrm{S}}(z)^{-1}=\dfrac{x_{r}T_{\text{r}}^{-1}+x_{\alpha}T_{\text{C}}^{-1}+x_{c}T_{\text{K}}^{-1}}{x_{r}+x_{\alpha}+x_{c}}$ (12)

### v2

@@ 2.2 Enhanced Pop II star-formation efficiency
Halos of mass $\sim 10^{5-10}\,\mathrm{M}_{\odot}$ are sensitive to the threshold for star formation $M_{\text{crit}}$, affected by feedback from streaming velocities, Lyman-Werner photons, and UV photons. The equations governing the threshold are as follows:
$M_{\text{vir}}=\dfrac{R_{\text{vir}}V_{c}^{2}}{G}\approx 5.8\times 10^{5}\mathrm{M}_{\odot}\left(\dfrac{V_{c}}{4.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{3}\left(\dfrac{1+z}{21}\right)^{-3/2};$ (3)
$M_{\text{mol}}\approx f_{v_{\text{bc}}}f_{\text{LW}}\left(5.8\times 10^{5}\mathrm{M}_{\odot}\right)\left(\dfrac{1+z}{21}\right)^{-3/2};$ (4)
$M_{\text{atm}}\approx\left(3.5\times 10^{7}\,\mathrm{M}_{\odot}\right)\left(\dfrac{1+z}{21}\right)^{-3/2}.$ (5)
Here, the first equation gives the virial mass of a halo with circular velocity $V_{c}$ (the minimum required for efficient cooling of gas in halos [^5] Note that this is simply related to a halo’s virial temperature often used in the literature via: $T_{\text{vir}}=\dfrac{\mu m_{\text{p}}V_{c}^{2}}{2k_{\text{B}}}\approx 635\,\text{K}\left(\dfrac{\mu}{0.6}\right)\left(\dfrac{V_{c}}{4.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{2}$ (6) where $m_{\text{p}}$ is the proton mass, $k_{\text{B}}$ is the Boltzmann constant, and $\mu$ is the mean molecular weight with $\mu=1.22$ for neutral primordial gas and $\mu=0.6$ for ionized hydrogen gas. We adopt the latter for consistency with Barkana & Loeb (2001) and other works, but the choice does not affect our results and simulations in any way as we directly deal with mass and velocity. It only affects our derived constraints on $T_{\text{vir}}$ presented in Table 3 and in Section 5.1.), the second equation is the molecular cooling threshold for halos (corresponding to $V_{c}\approx 4.2\,\mathrm{km}\,\mathrm{s}^{-1}$}), and the third is the atomic cooling threshold (corresponding to $V_{c}\approx 16.5\,\mathrm{km}\,\mathrm{s}^{-1}$}). The factors $f_{v_{\text{bc}}}$ and $f_{\text{LW}}$ account for the effects of streaming velocity and Lyman-Werner feedback on the molecular cooling threshold, respectively, as defined in Gessey-Jones et al. 2023 (see also Fialkov et al., 2012, 2013; Muñoz et al., 2022).
The variable $V_{c}$ encapsulates our uncertainty on the feedback processes in small halos (e.g. from SNe) and is a free parameter in our model, so that $M_{\text{crit}}(\mathbf{x},z)=\max\left(\min\left(M_{\text{mol}},M_{\text{atm}}\right),M_{\text{vir}}\right)$}. During the EoR, UV photons irradiate star-forming gas in ionized bubbles, further increasing $M_{\text{crit}}$ (to $\sim 10^{9}\,\mathrm{M}_{\odot}$) via photoheating feedback (described in Sobacchi & Mesinger, 2013; Cohen et al., 2016, accounting for ionized fractions in partially ionized cells).
@@
$f_{\star,\mathrm{II}}(\mathbf{x},M_{h},z)=\tilde{f}_{\star,\mathrm{II}}\begin{cases}0&M_{h}<M_{\text{crit}}\\ \dfrac{\log\left(M_{h}/M_{\text{crit}}\right)}{\log\left(M_{\text{atm}}/M_{\text{crit}}\right)}&M_{\text{crit}}\leq M_{h}<M_{\text{atm}}\\ 1&M_{\text{atm}}\leq M_{h}<M_{\text{high}}\\ \left(\dfrac{M_{h}}{M_{\text{high}}}\right)^{\alpha_{\star}}&M_{\text{high}}\leq M_{h}\end{cases}$ (7)
@@
and $f_{\star,\mathrm{II}}$ is naturally capped at $1$. The particular form of $M_{\text{high}}$ anchors the SFE at $z_{\text{min}}=6$, where our simulations end and the UVLF data is strongest (i.e., essentially calibrating the SFE magnitude and evolution to be with respect to redshift $z=6$). Figure 1 shows a schematic of this SFE model. Similar models with a flattening below a given mass threshold have been explored in the literature before (Sun & Furlanetto, 2016; Sipple & Lidz, 2024), the latter of whom find that a shallow-slope best fits the lensed UVLFs from HST at low halo masses, as opposed to a single power law (dash-dotted green in Figure 1) that assumes a steep decline. The parameterization used in our work, due to its flexibility, relaxes the assumption of declining SFE made by the single power law model for small mass halos. We also expect our model to be reasonable compared to the double power law model (solid pink line in Figure 1) for two reasons: (i) AGN feedback effects are generally expected for large halos $\sim 10^{12}\,\mathrm{M}_{\odot}$. However, few halos reach $\sim 10^{11}\,\mathrm{M}_{\odot}$ at $z\geq 6$ in our simulation volume of interest $\sim 10^{8}\,\mathrm{cMpc}^{3}$. For comparison, multi-field determinations of UVLFs combining many HST/JWST surveys typically probe volumes of $\lesssim 10^{6}\,\mathrm{cMpc}^{3}$ across redshift bins, making large halos even rarer. Since the 21-cm signal, and in general reionization, is sensitive to the cumulative effect of all halos down to $10^{5}\,\mathrm{M}_{\odot}$, our relevant halo mass range is wider than other abundance matching and analytic inferences that usually consider $M_{h}\gtrsim 10^{9}\,\mathrm{M}_{\odot}$. (ii) The role of AGN at high redshifts is still quite uncertain. Although pre-JWST studies suggest a decline in AGNs at $z\gtrsim 6$ (Kulkarni et al., 2019), more recent works indicate a non-negligible faint AGN population (Harikane et al., 2023b; Fujimoto et al., 2024). The contribution of these AGN to star-formation quenching remains unclear since the bright end of the luminosity function is poorly sampled at high redshifts.
The redshift evolution factor $\beta_{\star}$ is a timely addition, motivated by the abundance of bright galaxies in recent high-redshift JWST observations. Physically, the term captures an an enhancement in the SFE in the early Universe (for $\beta_{\star}>0$), meaning DM halos of the same mass host more massive and brighter galaxies in the early Universe than today.
@@ 2.2.1 Phenomenological comment on the SFE
The SFE defined here is the fraction of the accreted gas in a dark matter halo that is converted into stars, sometimes referred to as the integrated baryon conversion efficiency (e.g. Moster et al., 2018; Zhang et al., 2022). To first order, the SFE is equal to the stellar to halo mass fraction divided by the baryon fraction $f_{\text{b}}\approx 0.16$.
We note here that the SFE implementation in our sub-grid model is essentially an ensemble average since we do not track individual halos. For a given halo mass and redshift, the SFE of the galaxy in the host halo is fixed (i.e., there is no galaxy-to-galaxy scatter around the mean). This approach is reasonable as both the 21-cm signal and UVLF are statistical quantities probed at large scales (but see also Reis et al., 2022a; Nikolić et al., 2024, on the importance of stochasticity). Observations of individual galaxies can significantly differ from the ‘global mean’, and indeed any observational dataset is biased to the brightest and most massive galaxies which will be inherently stochastic (e.g Mirocha & Furlanetto, 2023). Nonetheless, rare galaxies manifest in our simulations via our pixel-level formalism. The abundance of halos and their gas-to-baryon mass fraction ($f_{\text{g}}/f_{\text{b}}$) is conditional on the halo mass, and pixel-level overdensity and streaming velocity. The minimum mass for star-forming halos $M_{\text{crit}}$ is further affected by local feedback mechanisms. Thus, the stellar mass and UVLF still varies between pixels across the simulation. Rare overdensities at high-$z$ will produce rare high mass galaxies, which can have an elevated SFE if they fall in the power-law regime of Equation 7. The main challenges of SFE enhancement (at any redshift) is feedback, and so there are two interpretations of this redshift evolution owing to differing star formation histories:
@@
    The first interpretation is one of bursty star-formation, where episodes of high-SFE can be induced by large-scale inflows (McClymont et al., 2025), positive feedback from AGN (Silk et al., 2024), or delayed stellar feedback (Furlanetto & Mirocha, 2022, which lead to a flattened SFE slope in the low halo mass regime; see their Figure 5). An extreme version of the last case are the “feedback-free bursts” (FFB) recently discussed in Dekel et al. (2023) and Li et al. (2024) where burstiness occurs on very short timescales of $\sim 10\,\mathrm{Myr}$ (i.e., multiple cycles/generations within a single half-crossing time of a virialized DM halo). This happens through rapid infall of gas via cold streams in low metallicity, high density environments, quicker than SNe/wind timescales, thereby essentially creating feedback-free conditions. The halo mass threshold above which these bursty galaxies form decreases with increasing redshift in this scenario, leading to an early time enchancement (see Figure 2 of Li et al., 2024).
@@
    The second interpretation is the standard uniform star-formation, through sustained gas supply and accretion to the galaxy over the halo half-crossing time. If the conditions in the early Universe are viable for compact galaxies hosting a large number of high density stellar clusters which have short dynamical timescales, they may be resilient to feedback processes essentially creating a “feedback failure” environment where stars can form with high efficiencies. These conditions could be due to increased binding energy (or deeper potential wells) of small halos at high-$z$ (Dayal et al., 2014; Furlanetto et al., 2017); dark-matter accelerated gas collapse (Boylan-Kolchin, 2025); or a SFE-scaling that increases with gas surface density analogous to giant molecular clouds (density modulated SFE or DMSFE; Somerville et al., 2025).
@@ 2.3.1 The 21-cm signal
where $n_{0}$ and $n_{1}$ are number densities of hydrogen atoms in the lower and higher energy states, the factor of $3$ accounts for the statistical degeneracy of the higher energy triplet state, $\nu_{21}=1420\,\mathrm{MHz}$} is the rest-frame frequency of the 21-cm line, and $T_{\mathrm{S}}$ is the spin temperature of the hydrogen atoms.
As the Universe expands, this 21-cm signal is observed today as a distortion to the CMB black body spectrum (or radio background, with temperature $T_{r}$) across the radio frequency range $\nu_{21,\text{obs}}=\nu_{21}/(1+z)$, corresponding to the different epochs at redshifts $z$. The strength of the signal is usually quantified in terms of the differential brightness temperature
$T_{21}(z)=\left(1-e^{-\tau_{21}}\right)\dfrac{T_{\mathrm{S}}-T_{r}}{1+z},$ (10)
@@
$\tau_{21}(z)\approx\dfrac{3}{32\pi}\dfrac{hc^{3}A_{10}}{k_{\text{B}}\nu_{21}^{2}}\dfrac{x_{\mathrm{HI}}(z)n_{\text{H}}(z)}{(1+z)dv_{\parallel}/dr_{\parallel}}\dfrac{1}{T_{\mathrm{S}}(z)},$ (11)
with $x_{\text{HI}}$ being the fractional abundance of neutral hydrogen, $n_{\text{H}}$ the hydrogen number density, $A_{10}$ the spontaneous emission rate of the 21-cm transition, and $dv_{\parallel}/dr_{\parallel}$ the proper velocity gradient along the observer line of sight. If $T_{\text{S}}>T_{r}$}, the signal is seen in net emission against the radio background, otherwise it is seen in net absorption.
In order to calculate this brightness temperature, one needs the spatially and temporally varying fractional abundance $x_{\mathrm{HI}}$ and the spin temperature $T_{\mathrm{S}}$. The former depends on the distribution of ionizing stellar/quasar sources, while the latter is an interplay between several competing influences that can cause a spin-flip transition in hydrogen. Coupling of $T_{\mathrm{S}}$ to the background radiation temperature $T_{r}$ is mediated via scattering of background photons, while coupling to the kinetic temperature of matter $T_{\text{K}}$ is through collisions between atoms/electrons in dense IGM, and the Wouthusen-Field (WF) effect when star-formation begins (Wouthuysen, 1952; Field, 1958). Thus, we get the equation (Furlanetto et al., 2006):
$T_{\mathrm{S}}(z)^{-1}=\dfrac{T_{r}^{-1}+x_{c}T_{\text{K}}^{-1}+x_{\alpha}T_{\text{C}}^{-1}}{1+x_{c}+x_{\alpha}}$ (12)